OpenStudy (anonymous):

Help!!!!!!!. Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = - x + 5. y = -2x + 2 y = 2x - 14 y = x - 8 y = -x - 4

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Is it y=-x-4

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (mathmale):

Hint: If two lines are parallel, their slopes are equal. I see only one answer choice that satisfies that.

OpenStudy (anonymous):

The correct answer is not one of those choices :)

OpenStudy (mathmale):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -x + 5. The slope is -1, and the given point is (4,-6). substituting these values into the point-slope form of the equation of a straight line, we get y-(-6)=-1(x-4). Try simplifying that.